OpenStudy (anonymous):

Why does 0.2 times 0.4 =0.08

OpenStudy (countryalltheway):

give me a sec

OpenStudy (countryalltheway):

because 0.4 times 2 equals 0.8, so times 0.2 would be 0.08

OpenStudy (zarkon):

\[0.2\times0.4=\frac{2}{10}\times\frac{4}{10}=\frac{2\times4}{10\times10}=\frac{8}{100}=0.08\]